# MLLMs Need 3D-Aware Representation Supervision for Scene Understanding

### v1

@@ 1 Introduction
In response to this challenge, recent state-of-the-art methods llava3d ; video3dllm ; scenellm ; 3d-llm ; gpt4scene have shifted towards leveraging multi-view images exclusively, drawing inspiration from the success of large-scale visual-language pretraining in multimodal LLMs (MLLMs) llava ; llava-onevision ; llavavideo ; gpt-4 ; qwen2vl ; gemini . These approaches aim to transfer 2D visual understanding to 3D scene comprehension by injecting 3D priors, such as 3D positional embeddings, into the models, thereby allowing MLLMs to capitalize on their extensive pretrained 2D knowledge for 3D interpretation.
@@
For the first question, we conduct comprehensive experiments to evaluate the 3D awareness on three representative MLLMs, including LLaVA-Next-Video llavavideo , LLaVA-One-Vision llava-onevision , and Qwen2-VL qwen2vl , following the finetuing settings of Video-3D LLM video3dllm . Specifically, we assess 3D awareness via view equivariance, quantifying it by computing the feature similarity between corresponding pairs from the same 3D voxel across different views. This evaluation requires MLLMs to associate the same object across multiple views, thereby reflecting their capacity for 3D representation. Our analysis encompasses five datasets spanning tasks such as 3D grounding scanrefer , captioning scan2cap , and question answering scanqa .
@@
In contrast, recent 3D foundation models such as VGGT vggt and FLARE flare are pretrained end-to-end on multi-view image sequences spanning a diverse set of 3D geometric tasks—including not only correspondence learning, but also depth estimation and camera parameter prediction. This comprehensive pretraining enables them to encode rich 3D properties within their features. Building on this, we propose a framework, 3DRS, that leverages these pretrained models by using their features as alignment targets for the visual outputs of MLLMs, thereby facilitating more effective 3D-aware representation learning. Unlike previous 3D MLLM approaches, in addition to traditional text token supervision, our framework employs explicit 3D-specific supervision directly on scene visual tokens. As demonstrated in our experiments (see Fig. 1), incorporating this form of 3D supervision consistently improves performance across a range of MLLMs and benchmarks. Notably, our approach incurs no additional training overhead, since the supervisory features can be pre-extracted offline. We believe this design offers valuable new insights for applying 3D foundation models in scene understanding. The key contribution of this paper can be summarized as follows:
@@ 2.1.1 Preliminaries
A MLLM typically consists of two main components: an image encoder $\mathcal{E}_{\mathrm{img}}$ and a text decoder $\mathcal{T}$. In this work, the input to our MLLM comprises a set of $N$ multi-view images =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{N}\}$, each associated with per-pixel 3D coordinates $\mathcal{C}=\{\mathbf{C}_{1},\mathbf{C}_{2},\ldots,\mathbf{C}_{N}\}$, where $\mathbf{C}_{i}\in\mathbb{R}^{H\times W\times 3}$ for image $I_{i}$ of size $H\times W$. The 3D coordinates for each pixel are computed from the depth map and the corresponding camera intrinsic and extrinsic parameters; detailed formulas and procedures can be found in the Sec. A.1.
@@ 2.1.2 Assessing 3D Feature Learning via Multi-View Correspondence
Inspired by the fact that cross-view correspondences are of crucial importance for 3D modeling multipleviewgeometric , we propose to use a correspondence-based evaluation framework. Given that the input consists of multi-view images, a natural metric for 3D feature quality is the strength of correspondence across views for the same 3D spatial location. This requires the model to associate and align objects or regions that occupy the same position in 3D space, regardless of viewpoint.
@@
To analyze the relationship between 3D feature learning and downstream task performance, we sort samples within each dataset by their correspondence scores and divide them into four quartiles (Q1–Q4, lowest to highest). From Fig. 2, we observe a clear trend: as the correspondence score increases, the model’s performance on the downstream task consistently improves. This strong positive correlation demonstrates the critical importance of 3D-aware representation quality for effective scene understanding in MLLMs.
@@ 2.2.1 Correspondence-based 3D Supervision Loss
By directly supervising positive pairs to be similar and negative pairs to be dissimilar, this correspondence loss encourages the model to learn multi-view 3D correspondences, thus enhancing the 3D-awareness of the learned representations. As will be shown in the experiments Sec. 3, supplementing the standard cross-entropy objective with $\mathcal{L}_{\mathrm{corr}}$ leads to improvements in downstream task performance. However, as this loss primarily targets view equivariance, the range of 3D properties captured remains limited, motivating the need for richer supervision.
@@
As shown in Fig. 3(a), given a set of multi-view images $\mathcal{I}$ for a scene, we first input them into a pretrained 3D foundation model $g$, which outputs a collection of per-pixel visual features $\{\mathbf{f}_{i}^{\mathrm{3D}}(p)\}$ for each image $I_{i}$ and pixel $p$. Since the spatial resolution of these features may differ from those of the MLLM outputs $\{\mathbf{f}_{i}(p)\}$, we apply 2D average pooling to the 3D foundation model’s output to match the MLLM feature map size.
@@
We then employ a cosine similarity-based distillation loss to maximize the alignment between the MLLM features $\hat{\mathbf{f}}_{i}(p)$ and the corresponding 3D foundation model features $\mathbf{f}_{i}^{\mathrm{3D}}(p)$:
@@
As illustrated in Fig. 3(b), we compare the correspondence scores before and after applying our 3DRS, where VGGT serves as the foundation model. The results consistently demonstrate that introducing 3DRS leads to substantial improvements in correspondence learning ability across all evaluated MLLMs and benchmarks. This proves the effectiveness of leveraging a pretrained 3D foundation model as a teacher model for enhancing 3D-aware representation learning in MLLMs. More comprehensive experimental results and analyses are detailed in Sec. 3.
@@ 3.1 Datasets and Evaluation Metrics
Datasets. We evaluate our approach on five benchmarks that collectively span key challenges in 3D scene understanding. ScanRefer scanrefer focuses on localizing objects using free-form language, while Multi3DRefer multi3drefer generalizes this to queries referencing zero, one, or multiple objects, better reflecting real-world ambiguity. Scan2Cap scan2cap addresses dense captioning by pairing detected objects in 3D scans with natural language descriptions. For question answering, ScanQA scanqa tasks models with answering open-ended questions grounded in 3D geometry and semantics, and SQA3D sqa3d goes further by requiring situated reasoning: agents must interpret their position and context to answer complex queries. All datasets are sourced from the richly annotated ScanNet scannet corpus, and we follow standard validation and test splits as established in prior work chatscene ; llava3d ; grounded-3dllm ; video3dllm . The statistics of training sets are detailed in the Sec. A.2.
@@
3.2 Implementation Details
3.3 Comparison with State-of-the-Art Models
Table 1 presents a comprehensive comparison between our approach, task-specific specialist models—which require fine-tuning on individual datasets—and 3D generalist models that are capable of handling multiple tasks. Compared to specialist models, our approach achieves substantial performance improvements. This demonstrates the significant benefits brought by joint training and the LLM-based architecture, which contribute to superior generalization and feature integration compared to methods tailored for specific tasks.
Furthermore, our method consistently outperforms 3D generalist approaches that utilize point clouds as input, such as LL3DA, Chat-3D, Grounded 3D-LLM, and 3D-LLaVA. Compared to Inst3D-LLM—which fuses multi-view images and point clouds—our approach also shows clear advantages, highlighting the strength of leveraging MLLMs as the backbone. Additionally, our method achieves considerable improvements over other MLLM-based methods, including LLaVA-3D and Video-3D LLM. These results collectively indicate that enhancing the 3D-awareness of MLLMs is highly effective for 3D scene understanding tasks, further validating the effectiveness of our proposed strategy.
@@ 3.4 Diagnostic Study
Effectiveness with different MLLMs. Table 2 demonstrates that integrating 3DRS with different MLLMs—LLaVA-Next-Video, LLaVA-OneVision, and Qwen2-VL—consistently boosts performance across all evaluated benchmarks. For example, LLaVA-Next-Video w/ 3DRS improves ScanRefer Acc@0.25 from 58.1 to 62.9, and Multi3DRef F1@0.25 from 58.0 to 60.4. Similar gains are observed for LLaVA-OneVision and Qwen2-VL, where 3DRS brings improvements on every dataset and metric. These results highlight the general applicability of our approach and its effectiveness in enhancing 3D scene understanding for various MLLMs.
Comparison between 2D and 3D foundation models. Table 3 compares the performance of using 2D and 3D foundation models as representation supervisors. It is clear that 3D foundation models (FLARE and VGGT) outperform all 2D foundation models (MAE, Siglip2, Dinov2) across almost every metric. This performance gap can be attributed to the inherent difference in the prior knowledge captured by 2D and 3D foundation models. 3D models are pre-trained on large-scale 3D data and thus better capture geometric structure, spatial relationships, and depth information, which are critical for 3D scene understanding tasks. In contrast, 2D foundation models, trained on images, lack explicit 3D spatial priors and struggle to provide effective supervision for learning 3D-aware representations. This highlights the importance of 3D-specific foundation models for achieving strong results in downstream 3D tasks.
Comparison of supervision signal. Table 5 shows that using correspondence loss for supervision brings improvements over the baseline, demonstrating the effectiveness of encouraging the model to learn multi-view correspondences. However, when 3D foundation model supervision is applied, the performance increases even further across all metrics. This indicates that 3D foundation models, with their rich 3D prior knowledge learned during pre-training, can more effectively enhance the 3D representation ability of MLLMs and yield greater gains for 3D understanding tasks.
Comparison of supervision at different layers. Table 5 examines the effect of applying 3D foundation model supervision at different layers of the network. The results reveal that supervision at deeper layers, especially the last layer, leads to the highest performance. This is likely because deeper layers are closer to the output and thus have a more direct impact on the final predictions. Additionally, these layers possess more parameters and a greater capacity to fit 3D features, which results in larger improvements on downstream tasks.
@@
Visualizations Fig. 4 illustrates qualitative results of our method across three tasks: visual grounding, object captioning, and question answering.
@@ 4.1 Scene Understanding with Large Language Models
4.2 3D-Awareness Study in Visual Models
Several studies have explored 3D-awareness; however, most of this research has focused on pure vision models rather than MLLMs, and primarily on 2D foundation models instead of 3D ones. For example, Probe3D probe3d and Lexicon3D lexicon3d conduct empirical analyses of the 3D-awareness in visual foundation models. CUA-O3D cua proposes integrating multiple 2D foundation models for 3D scene understanding, while Yang et al. multiview evaluates and enhances the 3D-awareness of ViT-based models for various downstream tasks.
In contrast to these works, our study centers on the 3D-awareness of MLLMs. Instead of enhancing 3D feature learning with 2D foundation models, we introduce 3D foundation models as supervisors to directly guide and improve the 3D representation ability of MLLMs.
@@ 5 Conclusion
In this paper, we systematically investigate the 3D representation capabilities of MLLMs for scene understanding. While previous studies have primarily focused on the impact of 2D foundation models, our work is the first to explore the value of 3D foundation models in this setting. Specifically, we introduce 3DRS, a framework that delivers direct 3D-aware supervision to MLLMs via pretrained 3D foundation models. Extensive experiments show that our approach consistently enhances performance across a range of 3D scene understanding tasks, underscoring the importance and potential of 3D foundation models for multimodal scene understanding.

### v2

# 3DRS: MLLMs Need 3D-Aware Representation Supervision for Scene Understanding

@@ 1 Introduction
In response to this challenge, recent state-of-the-art methods llava3d; video3dllm; scenellm; 3d-llm; gpt4scene have shifted towards leveraging multi-view images exclusively, drawing inspiration from the success of large-scale visual-language pretraining in multimodal LLMs (MLLMs) llava; llava-onevision; llavavideo; gpt-4; qwen2vl; team2023gemini. These approaches aim to transfer 2D visual understanding to 3D scene comprehension by injecting 3D priors, such as 3D positional embeddings, into the models, thereby allowing MLLMs to capitalize on their extensive pretrained 2D knowledge for 3D interpretation.
@@
For the first question, we conduct comprehensive experiments to evaluate the 3D awareness on three representative MLLMs, including LLaVA-Next-Video llavavideo, LLaVA-One-Vision llava-onevision, and Qwen2-VL qwen2vl, following the finetuing settings of Video-3D LLM video3dllm. Specifically, we assess 3D awareness via view equivariance, quantifying it by computing the feature similarity between corresponding pairs from the same 3D voxel across different views. This evaluation requires MLLMs to associate the same object across multiple views, thereby reflecting their capacity for 3D representation. Our analysis encompasses six datasets spanning tasks such as 3D grounding scanrefer, captioning scan2cap, and question answering scanqa.
@@
In contrast, recent 3D foundation models such as VGGT vggt and FLARE flare are pretrained end-to-end on multi-view image sequences spanning a diverse set of 3D geometric tasks—including not only correspondence learning, but also depth estimation and camera parameter prediction. This comprehensive pretraining enables them to encode rich 3D properties within their features. Building on this, we propose a framework, 3DRS, that leverages these pretrained models by using their features as alignment targets for the visual outputs of MLLMs, thereby facilitating more effective 3D-aware representation learning. Unlike previous 3D MLLM approaches, in addition to traditional text token supervision, our framework employs explicit 3D-specific supervision directly on scene visual tokens. As demonstrated in our experiments (see Fig.˜1), incorporating this form of 3D supervision consistently improves performance across a range of MLLMs and benchmarks. Notably, our approach incurs no additional training overhead, since the supervisory features can be pre-extracted offline. We believe this design offers valuable new insights for applying 3D foundation models in scene understanding. The key contribution of this paper can be summarized as follows:
@@ 2.1.1 Preliminaries
A MLLM typically consists of two main components: an image encoder $\mathcal{E}_{\mathrm{img}}$ and a text decoder $\mathcal{T}$. In this work, the input to our MLLM comprises a set of $N$ multi-view images $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{N}\}$, each associated with per-pixel 3D coordinates $\mathcal{C}=\{\mathbf{C}_{1},\mathbf{C}_{2},\ldots,\mathbf{C}_{N}\}$, where $\mathbf{C}_{i}\in\mathbb{R}^{H\times W\times 3}$ for image $I_{i}$ of size $H\times W$. The 3D coordinates for each pixel are computed from the depth map and the corresponding camera intrinsic and extrinsic parameters; detailed formulas and procedures can be found in the Sec.˜A.1.
@@ 2.1.2 Assessing 3D Feature Learning via Multi-View Correspondence
Inspired by the crucial role of cross-view correspondences in 3D modeling multipleviewgeometric, we propose a correspondence-based evaluation framework. Multi-view correspondences are fundamental in 3D vision, serving as essential cues for core tasks such as ray retriangulation multipleviewgeometric, bundle adjustment bundleadjustment, and pose estimation multiviewcategory. They are also critical for downstream applications like instance recognition and retrieval multiviewcategory; multishapereco; crossviewretrieval. Therefore, we adopt multi-view correspondence analysis as a proxy to evaluate the 3D representations of MLLMs. This approach requires the model to accurately associate and align objects or regions that occupy the same position in 3D space across different viewpoints.
@@
To analyze the relationship between 3D feature learning and downstream task performance, we sort samples within each dataset by their correspondence scores and divide them into four quartiles (Q1–Q4, lowest to highest). From Fig.˜2, we observe a clear trend: as the correspondence score increases, the model’s performance on the downstream task consistently improves. This strong positive correlation demonstrates the critical importance of 3D-aware representation quality for effective scene understanding in MLLMs.
@@ 2.2.1 Correspondence-based 3D Supervision Loss
By directly supervising positive pairs to be similar and negative pairs to be dissimilar, this correspondence loss encourages the model to learn multi-view 3D correspondences, thus enhancing the 3D-awareness of the learned representations. As will be shown in the experiments Sec.˜3, supplementing the standard cross-entropy objective with $\mathcal{L}_{\mathrm{corr}}$ leads to improvements in downstream task performance. However, as this loss primarily targets view equivariance, the range of 3D properties captured remains limited, motivating the need for richer supervision.
@@
As shown in Fig.˜3(a), given a set of multi-view images $\mathcal{I}$ for a scene, we first input them into a pretrained 3D foundation model $g$, which outputs a collection of per-pixel visual features $\{\mathbf{f}_{i}^{\mathrm{3D}}(p)\}$ for each image $I_{i}$ and pixel $p$. Since the spatial resolution of these features may differ from those of the MLLM outputs $\{\mathbf{f}_{i}(p)\}$, we apply 2D average pooling to the 3D foundation model’s output to match the MLLM feature map size.
@@
We then employ a distillation loss based on euclidean similarity to maximize the alignment between the MLLM features $\hat{\mathbf{f}}_{i}(p)$ and the corresponding 3D foundation model features $\mathbf{f}_{i}^{\mathrm{3D}}(p)$:
@@
As illustrated in Fig.˜3(b), we compare the correspondence scores before and after applying our 3DRS, where VGGT serves as the foundation model. The results consistently demonstrate that introducing 3DRS leads to substantial improvements in correspondence learning ability across all evaluated MLLMs and benchmarks. This proves the effectiveness of leveraging a pretrained 3D foundation model as a teacher model for enhancing 3D-aware representation learning in MLLMs. More comprehensive experimental results and analyses are detailed in Sec.˜3.
@@ 3.1 Datasets and Evaluation Metrics
Datasets. We evaluate our approach on six benchmarks that collectively span key challenges in 3D scene understanding. ScanRefer scanrefer focuses on localizing objects using free-form language, while Multi3DRefer multi3drefer generalizes this to queries referencing zero, one, or multiple objects, better reflecting real-world ambiguity. Scan2Cap scan2cap addresses dense captioning by pairing detected objects in 3D scans with natural language descriptions. For question answering, ScanQA scanqa tasks models with answering open-ended questions grounded in 3D geometry and semantics, and SQA3D sqa3d goes further by requiring situated reasoning: agents must interpret their position and context to answer complex queries. All these datasets are sourced from the richly annotated ScanNet scannet corpus, and we follow standard validation and test splits as established in prior work chatscene; llava3d; grounded-3dllm; video3dllm. Besides, VSI-Bench vsibench is used to evaluate the performance on visual-based spatial understanding tasks, which are composed of numerical and multiple-choice questions. The statistics of training sets are detailed in the Sec.˜A.2.
@@
3.2 Implementation Details
3.3 Comparison with State-of-the-Art Models
Table˜1 presents a comprehensive comparison between our approach, task-specific specialist models—which require fine-tuning on individual datasets—and 3D generalist models that are capable of handling multiple tasks. Compared to specialist models, our approach achieves substantial performance improvements. This demonstrates the significant benefits brought by joint training and the LLM-based architecture, which contribute to superior generalization and feature integration compared to methods tailored for specific tasks. Furthermore, our method consistently outperforms 3D generalist approaches that utilize point clouds as input, such as LL3DA, Chat-3D, Grounded 3D-LLM, and 3D-LLaVA. Compared to Inst3D-LLM—which fuses multi-view images and point clouds—our approach also shows clear advantages, highlighting the strength of leveraging MLLMs as the backbone. Additionally, our method achieves considerable improvements over other MLLM-based methods, including LLaVA-3D and Video-3D LLM. These results collectively indicate that enhancing the 3D-awareness of MLLMs is highly effective for 3D scene understanding tasks, further validating the effectiveness of our proposed strategy.
Table˜2 compares our method with leading proprietary APIs and open-source models on VSI-Bench vsibench, covering tasks such as object counting, spatial distance estimation, size reasoning, and sequence understanding. 3DRS achieves the best open-source results on most metrics—including object count, absolute distance, room size, and appearance oder—and remains competitive with proprietary models. These results demonstrate the strong spatial reasoning, generalization, and comprehensive scene understanding capabilities of our approach across diverse 3D vision tasks.
Table 2: Performance comparison on VSI-Bench.
Method Avg. Obj. Count Abs. Distance Obj. Size Room Size Rel. Dist. Rel. Dir. Route Plan Appr. Order
GPT-4o gpt-4 34.0 46.2 5.3 43.8 38.2 37.0 41.3 31.5 28.5
Gemini-1.5 Pro team2023gemini 45.4 56.2 30.9 64.1 43.6 51.3 46.3 36.0 34.6
LongVA-7B longva 29.2 38.0 22.2 33.1 43.3 25.4 15.7 33.1 17.7
InternVL2-40B internvl 37.0 41.3 26.2 48.2 27.5 47.6 32.7 27.8 44.7
LLaVA-Video-7B llavavideo 35.6 48.5 14.0 47.8 24.2 43.5 42.4 34.0 30.6
LLaVA-Video-72B llavavideo 40.9 48.9 22.8 57.4 35.3 42.4 36.7 35.0 48.6
LLaVA-OneVision-7B llava-onevision 32.4 47.7 20.7 47.4 42.5 35.2 29.4 24.4
LLaVA-OneVision-72B llava-onevision 40.2 43.5 23.9 57.6 37.5 42.5 39.9 32.5 44.6
3DRS 45.9 68.7 34.8 53.6 56.6 40.9 43.2 30.4 39.2
@@ 3.4 Diagnostic Study
Effectiveness with different MLLMs. Table˜3 demonstrates that integrating 3DRS with different MLLMs—LLaVA-Next-Video, LLaVA-OneVision, and Qwen2-VL—consistently boosts performance across all evaluated benchmarks. For example, LLaVA-Next-Video w/ 3DRS improves ScanRefer Acc@0.25 from 58.1 to 62.9, and Multi3DRef F1@0.25 from 58.0 to 60.4. Similar gains are observed for LLaVA-OneVision and Qwen2-VL, where 3DRS brings improvements on every dataset and metric. These results highlight the general applicability of our approach and its effectiveness in enhancing 3D scene understanding for various MLLMs.
Comparison between 2D and 3D foundation models. Table˜4 compares the performance of using 2D and 3D foundation models as representation supervisors. It is clear that 3D foundation models (FLARE and VGGT) outperform all 2D foundation models (MAE, Siglip2, Dinov2) across almost every metric. This performance gap can be attributed to the inherent difference in the prior knowledge captured by 2D and 3D foundation models. 3D models are pre-trained on large-scale 3D data and thus better capture geometric structure, spatial relationships, and depth information, which are critical for 3D scene understanding tasks. In contrast, 2D foundation models, trained on images, lack explicit 3D spatial priors and struggle to provide effective supervision for learning 3D-aware representations. This highlights the importance of 3D-specific foundation models for achieving strong results in downstream 3D tasks.
Comparison of supervision signal. Table˜6 shows that using correspondence loss for supervision brings improvements over the baseline, demonstrating the effectiveness of encouraging the model to learn multi-view correspondences. However, when 3D foundation model supervision is applied, the performance increases even further across all metrics. This indicates that 3D foundation models, with their rich 3D prior knowledge learned during pre-training, can more effectively enhance the 3D representation ability of MLLMs and yield greater gains for 3D understanding tasks.
Comparison of supervision at different layers. Table˜6 examines the effect of applying 3D foundation model supervision at different layers of the network. The results reveal that supervision at deeper layers, especially the last layer, leads to the highest performance. This is likely because deeper layers are closer to the output and thus have a more direct impact on the final predictions. Additionally, these layers possess more parameters and a greater capacity to fit 3D features, which results in larger improvements on downstream tasks.
@@ 4.1 Scene Understanding with Large Language Models
4.2 3D-Awareness in Vision Models
Several studies have investigated 3D-awareness; however, most prior work has focused on pure vision models rather than MLLMs, and primarily leveraged 2D foundation models instead of 3D ones. For example, FiT3D yue2024improving, Probe3D probe3d, and Lexicon3D lexicon3d empirically analyze the 3D-awareness of visual foundation models. CUA-O3D cua proposes integrating multiple 2D foundation models for 3D scene understanding, while Yang et al.multiview evaluates and enhances the 3D-awareness of ViT-based models in various downstream tasks. Some previous 3D detection works chen2022bevdistill; wang2023distillbev; bang2024radardistill have focused on improving 3D representations for pure vision models or CLIP-style vision-language models (VLMs), primarily aiming to enhance geometric understanding and spatial localization within unimodal or early multimodal frameworks. In addition, several studies on scene understanding 3d2ddis; pri3d; peng2023openscene have investigated various strategies for distilling 3D representations, such as transferring knowledge from 3D models to 2D networks or promoting cross-modal alignment. However, these efforts have not addressed the unique challenges presented by MLLMs, which require a more holistic integration of visual, linguistic, and spatial information. As a result, the potential of distilling 3D awareness into MLLMs for richer and more generalizable scene understanding remains largely unexplored.
In contrast, our work specifically targets the 3D-awareness of MLLMs. Rather than enhancing 3D feature learning via 2D foundation models, we introduce 3D foundation models as supervisors to directly guide and improve the 3D representation capabilities of MLLMs.
@@ 5 Conclusion
In this paper, we present a comprehensive study of the 3D representation capabilities of multi-modal large language models (MLLMs) in the context of scene understanding. While most existing research has centered on leveraging 2D foundation models to improve visual reasoning in MLLMs, the role and utility of 3D foundation models in this setting remain largely unexplored. To bridge this gap, we propose 3DRS, a novel framework that introduces direct 3D-aware supervision to MLLMs by leveraging pretrained 3D foundation models as teachers. Our approach enables MLLMs to acquire richer geometric and spatial representations, facilitating more accurate and robust understanding of complex 3D scenes. Through extensive experiments on diverse 3D scene understanding benchmarks, we demonstrate that 3DRS consistently improves performance across a variety of tasks, such as object localization, spatial reasoning, and 3D question answering. These results highlight the unique advantages and significant potential of integrating 3D foundation models for advancing multimodal scene understanding.
6 Limitation
While our paper aims to enhance the 3D-awareness of MLLMs, the relatively limited size of the dataset used for finetuning—especially when compared to that used during the MLLM pretraining stage—may restrict the full realization of our approach’s potential. Consequently, the improvements demonstrated in this work may only represent an initial step toward more robust 3D understanding. A promising direction for future research is to incorporate 3D-awareness learning into the pretraining stage of MLLMs, which could lead to fundamentally stronger models with deeper 3D comprehension. Besides, due to the distillation-based nature of our approach, the performance of our method is upper-bounded by the quality of the teacher 3D foundation model. Any limitations or failure modes of the teacher—such as inaccurate correspondence, erroneous depth estimation, or incomplete geometric representations—can be propagated to the student MLLM and may potentially mislead it during training. While our experiments demonstrate consistent improvements over strong baselines, it is possible that errors or biases in the teacher’s predictions can negatively impact the downstream 3D reasoning abilities of the student model. We believe that, as 3D foundation models continue to rapidly advance, this limitation becomes less pronounced over time.
Acknowledgments
This work is supported by Hong Kong Research Grant Council - General Research Fund (Grant No. 17213825). We would like to thank Weining Ren for the valuable and insightful discussions.